OpenStudy (anonymous):

Photocopying. Quick Copy recently charged 49¢ per page for color copies and 7¢ per page for black-and-white copies. If Shirlee’s bill for 90 copies was $11.34, how many copies of each type were made?

OpenStudy (anonymous):

49c+7b=11.34*100. c+b=90 solve these two equation and get the value of c and b. where c=no of color copies and b= no of black n white copies

OpenStudy (cwrw238):

let number of black and white copies = x and pf color = y 49y + 7x = 1134 x + y = 90

OpenStudy (anonymous):

Ok I have that part figured out I am stuck after that

OpenStudy (anonymous):

Then use substitution by substituting 90-c for b. Your new equation would then be: 49c + 7(90 - c) = 11.34*100

OpenStudy (cwrw238):

to solve x secong equation by 7 7x + 7y = 630 now subtract this from ist equaton to eliminate x: 42y = 504 so y = 12 and x = 90 - 7 = 83 so there are 83 copies of black af collor

OpenStudy (cwrw238):

83 of black/white and 7 of color

OpenStudy (anonymous):

ok thanks

OpenStudy (cwrw238):

soorry = typo there 83 of b/w and 12 of color ( not 7!)

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

shhoot - i'm an idiot its 12 color and 90 - 12 = 78 b/w check it out : - 12*49 + 78* 7 = 1134 - yess!! correct answer 12 color 78 b/w

OpenStudy (cwrw238):

sorry - the grand kids are making a lot of noise tonight!!!

OpenStudy (anonymous):

that ok i understand

OpenStudy (cwrw238):

lol - got there in the end!

OpenStudy (anonymous):

thanks again